# Probabilistically Tightened Linear Relaxation-based Perturbation Analysis for Neural Network Verification

### v1

@@ Abstract
We present Probabilistically Tightened Linear Relaxation-based Perturbation Analysis (PT-LiRPA), a novel framework that combines over-approximation techniques from LiRPA-based approaches with a sampling-based method to compute tight intermediate reachable sets. In detail, we show that with negligible computational overhead, PT-LiRPA exploiting the estimated reachable sets, significantly tightens the lower and upper linear bounds of a neural network’s output, reducing the computational cost of formal verification tools while providing probabilistic guarantees on verification soundness. Extensive experiments on standard formal verification benchmarks, including the International Verification of Neural Networks Competition, show that our PT-LiRPA-based verifier improves robustness certificates by up to 3.31X and 2.26X compared to related work. Importantly, our probabilistic approach results in a valuable solution for challenging competition entries where state-of-the-art formal verification methods fail, allowing us to provide answers with high confidence (i.e., at least 99%).
@@ 1 Introduction
Deep neural networks (DNNs) and recently large language models (LLMs) have revolutionized various fields, from healthcare and finance to natural language processing, enabling remarkable capabilities, for instance, in image recognition [33], robotic tasks such as manipulation [36, 27] and navigation [41, 28, 13]. Nonetheless, their opacity and vulnerability to the so-called “adversarial inputs" [40, 1] have also raised significant concerns, especially when deployed in safety-critical applications such as autonomous driving or medical diagnosis. Consequently, developing methods to certify the safety aspect of these models is crucial. Achieving provable safety guarantees involves employing formal verification (FV) of neural network techniques [23], which mathematically ensures that a system will never produce undesired outcomes in all configurations tested.
In this work, we focus on robustness verification, which aims to guarantee that a model’s output remains consistently robust under small predefined input perturbations around a given point ${\bm{x}}_{0},$ typically defined as an $\ell_{\infty}$ ball of $\varepsilon$ radius around the input ${\bm{x}}_{0},$ i.e., the set $\mathcal{C}=\mathcal{C}_{{\bm{x}}_{0},\epsilon}=\{{\bm{x}}|\;||{\bm{x}}-{\bm{x}}_{0}||_{\infty}\leq\epsilon\}.$ A standard approach to proving the robustness property is to first encode the desired output of the classifier as the sign of the value $f$ computed by a single (output) node, i.e., so that the correct classification in ${\bm{x}}_{0}$ corresponds to $f({\bm{x}}_{0})>0.$ Then, the robustness verification becomes equivalent to checking that $\min\limits_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})$, is positive [44]. However, due to the non-linear and non-convex nature of the DNN, solving this problem is, in general, NP-hard [22, 47]. A recent line of works called linear relaxation-based perturbation analysis (LiRPA) algorithms [54, 52, 44, 51] proposes a formal analysis based on a sound linear relaxation of the DNN. The idea is to compute relaxations of all sources of non-linearity in the network so as to obtain two linear functions providing respectively a lower and an upper bound of the DNN’s output, which are then used in the robustness certification. Nonetheless, these methods use overapproximation techniques to satisfy the worst-case setting, where no exceptions are allowed. While this strict approach ensures absolute safety within the certified input space, it faces crucial limitations. In fact, the efficiency of LiRPA approaches, and, in general, any FV methods, scales poorly with the size of the model. Additionally, formal methods fail to offer meaningful robustness information when input perturbations exceed the certified bounds. As discussed in [48], relying solely on formal verification can lead to overly conservative outcomes, particularly when adversarial examples are rare or when prior knowledge of the input distribution is available. This stems from the exact nature of formal solvers, which treat all violations equally: even a single adversarial input—such as a one-pixel change—invalidates the entire region, without distinguishing between isolated, low-probability cases and large, semantically significant unsafe regions.
Similar to recent approaches [48, 12], to overcome these critical limitations, we propose a probabilistic perspective that allows for an infinitesimal trade-off in certainty while providing the likelihood of such violations, enabling more nuanced and practical assessments of robustness that better reflect real-world risks. Importantly, the goal of this work is not to compare probabilistic methods with provable ones, as they are fundamentally different in nature. Rather, our objective is to propose a novel, complementary solution that enhances formal methods by offering additional safety information about the models under evaluation for particular challenging instances to be verified.
@@
Building on this foundation, we explore Xu et al. speculation that estimates as tightly as possible reachable sets, i.e., the range of possible values at each hidden node given bounded inputs, could significantly enhance verification efficiency by yielding sharper final linear bounds. In this perspective, we address two key research questions: (i) How can we compute reachable sets trading off (probabilistically measurable) tightness and computational tractability? (ii) What theoretical guarantees can we establish for linearized layers using (probabilistically sound) reachable set?
Our Contributions.
@@
We propose Probabilistically Tightened Linear Relaxation-based Perturbation Analysis (PT-LiRPA), a novel framework for computing tight linear lower and upper bounds combining existing LiRPA methods with a sampling-based reachable set estimation strategy. Unlike other related probabilistic works [48, 12], our approach does not rely on specific input distributions or attacks. Specifically, in this work, we start considering statistical results on tolerance limits [50] to provide probabilistic guarantees on estimating the minimum and maximum values of a neural network’s output. Using the approach of Wilks, we are guaranteed that, for any $R,\psi\in(0,1),$ with probability $\psi$, at most, a fraction $(1-R)$ of points from a possibly infinitely large sample in the perturbation region $\mathcal{C}$ may violate the estimated bounds obtained from an initial sample, whose size $n$ is explicitly computable from the desired parameters $\psi$ and $R$. Hence, our intuition is to extend this approach to also compute an estimation of reachable sets with a specified confidence level. Nonetheless, while the result of [50] quantitatively bounds the error of the sample-based procedure by predicting the fraction of potential violations in future samples, it results in "weaker" guarantees with respect to other known probabilistic approaches. In particular, this approach does not provide information on the magnitude of these violations with respect to the estimated bounds. Specifically, in any estimated reachable set for a possibly existing fraction $(1-R)$ of points drawn from $\mathcal{C}$, the probability of violating the bound might in principle be uncontrollably large. To address such an issue, following results on extreme value theory (EVT) [19, 15], we extend Wilks [50]’ probabilistic guarantees and present two novel bounds on the magnitude of potential errors between the true minimum (Theorem 3.5) and the sample-based estimated one (Theorem 3.7). Notably, our final result proves that with negligible computational overhead, each node’s reachable set computed using random samples represents with high probability the actual domain of that node for any $x\in\mathcal{C}$. Hence, we show that by employing this sampling-based procedure to compute probabilistically tight reachable set bounds in the neural network and integrating these into the linearization used by LiRPA-based formal verification methods, we are able to obtain significantly tighter lower and upper linear bounds of a neural network’s output, while preserving the verification soundness for any input in the perturbation region and specified confidence level.
@@
To assess the benefit and effectiveness of our novel framework, we perform an extensive empirical evaluation. We first compare our approach on neural networks trained on MNIST and CIFAR datasets with PROVEN [48] and Randomized Smoothing [12], the most closely related probabilistic verification approaches. In addition, as a ground truth, we also consider a set of state-of-the-art worst-case robustness verification approaches, namely CROWN[54], $\alpha$-CROWN [52], $\beta$-CROWN [44], and GCP-CROWN [55]. This first set of experiments shows that with a very high confidence (i.e., $\geq 99\%$), PT-LiRPA improves the certified lower bound of $\varepsilon$ perturbation tolerated by the models up to 3.31x and 2.26x compared with both the corresponding probabilistic approaches of [48, 12] and up to 3.62x w.r.t. the worst-case analysis results. Finally, in the second set of experiments, we demonstrate that for challenging instances from the International Verification of Neural Networks Competition (VNN-COMP) [32, 8], where state-of-the-art formal verification methods fail to produce a conclusive result, our PT-LiRPA, with a quantifiable level of confidence represents a valuable resource in providing safety information.
@@
2 Preliminaries and Related Work
@@ 2.1 Notation and Problem Formulation
As observed in the introduction, we assume, without loss of generality, that there is a single node in the $N$th layer, which we simply denote by $z^{(N)}$.111 One can simply enforce this condition for networks that do not satisfy this assumption by adding one layer and encoding, for instance, the robustness property that we aim to verify in a single output node as a margin between logits, which produces a positive output only if the correct label is predicted [23, 44]. Hence we have $f({\bm{x}})=\hat{{\bm{z}}}^{(N)}({\bm{x}})={z}^{(N)}({\bm{x}}).$
In our following arguments, we assume that the activation function of each node is the ReLU, which is the most employed in the verification works of the literature [52, 44], but the soundness of the proposed approach can also be extended for different non-linear scalar functions studied in the literature, such as Tanh, Sigmoid, GeLU, as in [51].
Hence, we define the robustness verification problem of deep neural networks as follows.
Given an input point of interest ${\bm{x}}_{0}$, for which $f({\bm{x}}_{0})>0$, and an input perturbation set $\mathcal{C}=\mathcal{C}_{{\bm{x}}_{0},\epsilon}=\{{\bm{x}}|\;||{\bm{x}}-{\bm{x}}_{0}||_{\infty}\leq\epsilon\}$, i.e., $\mathcal{C}$ is an $\ell_{\infty}$ ball of $\varepsilon$ radius around an original input ${\bm{x}}_{0}$, we aim to find, if there exists, an input ${\bm{x}}\in\mathcal{C}$ such that $f({\bm{x}})\leq 0$, thus resulting in a violation of the property. If $f({\bm{x}})>0\;\forall{\bm{x}}\in\mathcal{C}$, we say $f$ is robust (or verified) to all the possible input perturbations in $\mathcal{C}$. A possible way to prove the property is to solve the optimization problem in terms of $\min\limits_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})$ and by checking if the result is positive. Formally:
@@
Because of the effect of the activation functions $\sigma$ applied to the value computed in each node, the resulting function $f$ computed by the network is, in general, non-convex, making the above Robustness verification problem NP-hard [22]. In order to cope with this issue, (in)complete verifiers usually relax the DNNs’ non-convexity to obtain over-approximate sound lower and upper bounding functions on $f$, respectively, denoted by $\underline{f}$ and $\overline{f},$ i.e., $\underline{f}({\bm{x}})\leq f({\bm{x}})\leq\overline{f}({\bm{x}})$ for all ${\bm{x}}\in\mathcal{C}.$ Therefore, if $\underline{f}^{*}=\min_{{\bm{x}}\in\mathcal{C}}\underline{f}({\bm{x}})>0$, then also $f^{*}=\min_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})>0$, i.e., the real minimum value of $f$ will be positive, and similarly if $\overline{f}^{*}=\min_{{\bm{x}}\in\mathcal{C}}\overline{f}({\bm{x}})\leq 0$ than also $f^{*}=\min_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})\leq 0$.
In both these situations, we can return a provable result. However, if $\underline{f}^{*}<0<\overline{f}^{*}$, we cannot be sure about the sign of $f^{*}$, and we typically have to proceed with a branch and bound (BaB) [10] process. More specifically, many FV tools firstly recursively divide the original verification problem into smaller subdomains, either by dividing the perturbation region [43] or by splitting ReLU neurons into positive/negative linear domains [9]. Secondly, they bound each subdomain with specialized (incomplete) verifiers, typically linear programming (LP) solvers [17], which can fully encode neuron split constraints. The verification process ends once we verify all the subdomains of this searching tree, or we find a single counterexample ${\bm{x}}$ such that $\overline{f}({\bm{x}})\leq 0$. Even though LP-verifiers are mainly used in complete FV tools, recent LiRPA-based approaches [52, 44] show how to solve an optimization problem that is equivalent to the costly LP-based methods with neuron split constraints while maintaining the efficiency of bound propagation techniques, significantly outperforming LP-verification time thanks to GPU acceleration.
LiRPA approaches [54, 38, 51, 52, 44] propose to cope with the non-linearity of the function computed by a neural network by computing linear approximations of each non-linear unit. The high-level idea is to compute bounds on each neuron’s function in the DNN, for instance, all the ReLU nodes, that can be expressed by linear functions, for which the above robustness verification problem can be efficiently solved. In detail, using interval bound propagation (IBP) [24], we first compute a reachable set for each neuron $z^{(i)}_{j}$, defined as the interval $[l^{(i)}_{j},u^{(i)}_{j}]$ of its pre-activation values, such that $l^{(i)}_{j}\leq z^{(i)}_{j}({\bm{x}})\leq u^{(i)}_{j}$, for all ${\bm{x}}\in\mathcal{C}.$ The ReLU post-activation value of the node is given by $\hat{z}^{(i)}_{j}=\max(0,z^{(i)}_{j}$. Therefore, the node is considered “unstable" if its pre-activated bounds are such that $u^{(i)}_{j}>0>l^{(i)}_{j}$ and a linear approximate bound can be computed as depicted in Fig. 1.
Figure 1: Linear relaxation for ReLU($z^{(i)}_{j}$)
@@ 2.2 Linear Relaxation-based Perturbation Analysis (LiRPA) approaches
Otherwise, the node is either considered “active" if $l^{(i)}_{j}\geq 0$ or “inactive" if $u^{(i)}_{j}\leq 0$. Once linear bounds are established across all neurons, two propagation methods are typically employed: forward and backward. In forward propagation, the linear bounds for each neuron are expressed in terms of the input and propagated layer by layer until the output is reached. In backward propagation, we start from the output and propagate the bounds backward to earlier layers until we can express a linear relation between input and output.
@@
In detail, to improve the tightness of the bounds, [51] proposes a refined backward computation strategy that leverages information from a prior forward propagation. After computing all reachable sets using a method such as interval bound propagation [24], the backward analysis constructs layer-wise linear relaxations starting from the output layer back to the input.
For clarity purposes, in the following we will only report the notation for the linear lower bound computation, but similar considerations also apply to the upper bound.

### v2

# Luca Marzari, Ferdinando Cicalese and Alessandro Farinelli

@@ Abstract
We present Probabilistically Tightened Linear Relaxation-based Perturbation Analysis (PT-LiRPA), a novel framework that combines over-approximation techniques from LiRPA-based approaches with a sampling-based method to compute tight intermediate reachable sets. In detail, we show that with negligible computational overhead, PT-LiRPA exploiting the estimated reachable sets, significantly tightens the lower and upper linear bounds of a neural network’s output, reducing the computational cost of formal verification tools while providing probabilistic guarantees on verification soundness. Extensive experiments on standard formal verification benchmarks, including the International Verification of Neural Networks Competition, show that our PT-LiRPA-based verifier improves robustness certificates, i.e., the certified lower bound of $\varepsilon$ perturbation tolerated by the models, by up to 3.31X and 2.26X compared to related work. Importantly, our probabilistic approach results in a valuable solution for challenging competition entries where state-of-the-art formal verification methods fail, allowing us to provide answers with high confidence (i.e., at least 99%).
@@ 1 Introduction
Deep neural networks (DNNs) and recently large language models (LLMs) (vaswani2017attention) have revolutionized various fields, from healthcare and finance to natural language processing, enabling remarkable capabilities, for instance, in image recognition (image), robotic tasks such as manipulation (manipulation1; manipulation2) and navigation curriculum; navigation; TIST; eps_retrain. Nonetheless, their opacity and vulnerability to the so-called “adversarial inputs" (adversarial; amir2022verifying) have also raised significant concerns, especially when deployed in safety-critical applications such as autonomous driving or medical diagnosis. Consequently, developing methods to certify the safety aspect of these models is crucial. Achieving provable safety guarantees involves employing formal verification (FV) of neural network techniques (LiuSurvey), which mathematically ensures that a system will never produce undesired outcomes in all configurations tested.
[…]
@@
Building on this foundation, we explore acrown speculation that estimates as tightly as possible reachable sets, i.e., the range of possible values at each hidden node given bounded inputs, could significantly enhance verification efficiency by yielding sharper final linear bounds. In this perspective, we address two key research questions: (i) How can we compute (probabilistic) reachable sets that yield tighter intermediate and output bounds than existing methods, while remaining computationally efficient? (ii) What theoretical guarantees can we establish for linearized layers using (probabilistically sound) reachable set?
@@
Our Contributions. We propose Probabilistically Tightened Linear Relaxation-based Perturbation Analysis (PT-LiRPA), a novel framework for computing tight linear lower and upper bounds combining existing LiRPA methods with a sampling-based reachable set estimation strategy. Unlike other related probabilistic works (proven; randomizedSmoothing), our approach does not rely on specific input distributions or attacks. Specifically, in this work, we start considering statistical results on tolerance limits (wilks) to provide probabilistic guarantees on estimating the minimum and maximum values of a neural network’s output. Using the approach of wilks, we are guaranteed that, for any $R,\psi\in(0,1),$ with probability $\psi$, at most, a fraction (1R)(1-R) of points from a possibly infinitely large sample in the perturbation region $\mathcal{C}$ may violate the estimated bounds obtained from an initial sample, whose size $n$ is explicitly computable from the desired parameters $\psi$ and $R$. Hence, our intuition is to extend this approach to also compute an estimation of reachable sets with a specified confidence level. Nonetheless, while the result of (wilks) quantitatively bounds the error of the sample-based procedure by predicting the fraction of potential violations in future samples, it results in "weaker" guarantees with respect to other known probabilistic approaches. In particular, this approach does not provide information on the magnitude of these violations with respect to the estimated bounds. Specifically, in any estimated reachable set for a possibly existing fraction (1R)(1-R) of points drawn from $\mathcal{C}$, the probability of violating the bound might in principle be uncontrollably large. To address such an issue, following results on extreme value theory (EVT) (haan2006extreme; de1981estimation), we extend wilks’ probabilistic guarantees and present two novel bounds on the magnitude of potential errors between the true minimum (Theorem 3.5) and the sample-based estimated one (Theorem 3.7). Notably, our final result proves that with negligible computational overhead, each node’s reachable set computed using random samples represents with high probability the actual domain of that node for any $x\in\mathcal{C}$. Hence, we show that by employing this sampling-based procedure to compute probabilistically tight reachable set bounds in the neural network and integrating these into the linearization used by LiRPA-based formal verification methods, we are able to obtain significantly tighter lower and upper linear bounds of a neural network’s output, while preserving the verification soundness for any input in the perturbation region and specified confidence level.
@@
To assess the benefit and effectiveness of our novel framework, we perform an extensive empirical evaluation. We first compare our approach on neural networks trained on MNIST and CIFAR datasets with PROVEN (proven) and Randomized Smoothing (randomizedSmoothing), the most closely related probabilistic verification approaches. In addition, as a ground truth, we also consider a set of state-of-the-art worst-case robustness verification approaches, namely CROWN(crown), $\alpha$-CROWN (acrown), $\beta$-CROWN (bcrown), and GCP-CROWN (GCP-CROWN). This first set of experiments shows that with a very high confidence (i.e., $\geq 99\%$), PT-LiRPA improves the certified lower bound of $\varepsilon$ perturbation tolerated by the models up to 3.31x and 2.26x compared with both the corresponding probabilistic approaches of (proven; randomizedSmoothing) and up to 3.62x w.r.t. the worst-case analysis results. Finally, in the second set of experiments, we demonstrate that for challenging instances from the International Verification of Neural Networks Competition (VNN-COMP) (VNN-comp2022; VNN-comp2023), where state-of-the-art formal verification methods fail to produce a conclusive result, our PT-LiRPA, with a quantifiable level of confidence represents a valuable resource in providing safety information.
@@
2 Related Work and Preliminaries
2.1 Related Work
Formal Verification.
In recent years, significant research has been dedicated to formal verification and especially to robustness verification (LiuSurvey; wei2024modelverification). For example, (jair1) shares with our work the goal of achieving certified robustness for neural networks, but differs significantly in both focus and methodology. Their approach is specifically designed for the time-series domain, leveraging statistical constraints and polynomial transformations to generate adversarial examples and derive robustness guarantees. In contrast, our work addresses general classification tasks and introduces a probabilistically sound verification method based on tight linear relaxation bounds. Other related work that integrates recurrent neural network(RNN)-based policy learning with formal verification is presented in (jair2). Specifically, the authors target policy verification under temporal logic constraints in partially observable setups by extracting finite-state controllers from RNNs to enable model checking. In contrast, our work focuses on classification tasks and specifically robustness verification of general feedforward deep neural networks under input perturbations. For this type of verification is important to cite sound and complete verifiers such as mixed integer programming (MIP) MIP and satisfiability module theory(SMT)-based solvers Reluplex; marabou2. The most closely related works to our proposal comprise LiRPA-based verification approaches that focus on increasing the quality of linear bounds of the most popular activation functions, such as ReLU, and more general activation functions. More specifically, (autolirpa) proposes a framework for deriving and computing near-optimal sound bounds with linear relaxation-based perturbation analysis for neural networks. This framework is the base of all the most famous state-of-the-art formal verification tools such as CROWN (crown), $\alpha$-CROWN (acrown), $\beta$-CROWN (bcrown), GCP-CROWN (GCP-CROWN), the top performer on last years VNN-COMP (VNN-comp2022; VNN-comp2023). Notably, this approach has also been recently employed to both provably (kotha2023provably) and approximate (zhang2024provable) bounding neural network preimages.
All these approaches will be used as a worst-case verification result in our experiments.
Different approaches have tried to incorporate a sampling-based approach to enhance either the linear relaxation of arbitrary non-linear functions (linsyn; sol; dualapp) or the verification process (geometric_sample), but still maintaining provable verification certificates. For instance, (sol; linsyn) proposed a method synthesizing linear bounds for arbitrary complex activation functions, such as GeLU (gelu) and Swish (swish), by combining a sampling technique with an LP solver to synthesize candidate lower and upper bound coefficients and then certifying the final result via SMT solvers (smtsolver). In (dualapp), the authors focus on estimating the actual domain of an activation function by combining Monte Carlo simulation and gradient descent methods to compute an underestimated domain, which is then paired with over-approximations to define provable linear bounds. If, on the one hand, our PT-LiRPA also employs a similar sampling-based procedure to compute the estimated domain of reachable sets, on the other hand, it differs fundamentally in both nature and focus. Specifically, our approach provides an explicit formula that, for any desired level of confidence, gives the number of samples sufficient to estimate the minimum and maximum pre-activation value of any node (over the set of inputs in $\mathcal{C}$) with the given confidence. In addition, we are also able to estimate the maximum error between our estimates and the true extremal pre-activation values. Crucially, our approach adopts a probabilistic perspective to derive safety insights, whereas (dualapp) relies on combining under- and over-approximations to produce provable bounds—an approach that may inherit the limitations of formal methods discussed earlier.
Probabilistic Verification.
Recently, several works have explored probabilistic verification of machine learning models. For example, (probabilistic1; probabilistic2; probabilistic9) focus on a probabilistic verification perspective for general machine learning models (e.g., deep generative/diffusion models), leveraging either uncertainty sources generated by specific encoders or using symbolic reasoning and probabilistic inference. (probabilistic3) proposes a Monte Carlo-based method utilizing multi-level splitting to estimate the probability of rare events for robustness verification. Other approaches rely on Chernoff-Cramér bounds, e.g., (probabilistic8) estimates the local variation of neural network mappings at training points to regularize the loss function. In (probabilistic10), the authors probabilistically certify a neural network by overapproximating input regions where robustness is violated. In (probabilistic4) a probabilistic certification approach is proposed that can be used in general attack settings to estimate the probability of a model failing if the attack is sampled from a certain distribution. Another line of work, like (CLEVER) uses extreme value theory (EVT) (haan2006extreme) to statistically estimate the local Lipschitz constant and assess the probabilistic robustness of the model without relying on the overapproximation of LiRPA-based solvers. Finally, recent works (probabilistic5; marzari2023dnn; marzari2024enumerating; probabilistic6; probabilistic7) focus either on different types of verification, such as probabilistic enumeration of (un)safe region for neural networks, or on probabilistic specifications for robustness verification, which falls outside the scope of this paper, as we are interested in standard robustness verification with common specifications.
More closely related to our work are the approaches used in PROVEN (proven) and Randomized Smoothing (randomizedSmoothing), which also focus on probabilistic robustness guarantees for neural network classifiers. PROVEN builds upon LiRPA-based techniques, combining them with concentration inequalities to derive probabilistic bounds on the network’s output under input perturbations. Similarly, Randomized Smoothing constructs a smoothed classifier by adding random noise, typically Gaussian, to the input and then provides robustness guarantees by analyzing the output distribution of the smoothed model. PT-LiRPA aligns with these approaches in that it also leverages linear relaxation techniques and probabilistic reasoning, but introduces a novel perspective by focusing on how much sampling-based underestimations affect the intermediate linear bounds used during network linearization. This enables a finer-grained analysis of robustness with strong probabilistic guarantees, which is particularly useful in scenarios where the available methods for computing exact worst-case bounds either fail (by not terminating under reasonable time and space constraints) or terminate with only over-conservative bounds.
@@ 2.2 Notation and Problem Formulation
As observed in the introduction, we assume, without loss of generality, that there is a single node in the $N$th layer, which we simply denote by $z^{(N)}$. [^1] One can simply enforce this condition for networks that do not satisfy this assumption by adding one layer and encoding, for instance, the robustness property that we aim to verify in a single output node as a margin between logits, which produces a positive output only if the correct label is predicted (LiuSurvey; bcrown). Hence we have $f({\bm{x}})=\hat{{\bm{z}}}^{(N)}({\bm{x}})={z}^{(N)}({\bm{x}}).$
In our following arguments, we assume that the activation function of each node is the ReLU, which is the most employed in the verification works of the literature (acrown; bcrown), but the soundness of the proposed approach can also be extended for different non-linear scalar functions studied in the literature, such as Tanh, Sigmoid, GeLU, as in (autolirpa). Hence, we define the robustness verification problem of deep neural networks as follows.
[…]
@@
Output: $\texttt{Robust}\iff$ $\min\limits_{{\bm{x}}\in\mathcal{C}}f({\bm{x}}):=z^{(N)}({\bm{x}})>0$.
Because of the effect of the activation functions $\sigma$ applied to the value computed in each node, the resulting function $f$ computed by the network is, in general, non-convex, making the above Robustness verification problem NP-hard (Reluplex). In order to cope with this issue, (in)complete verifiers usually relax the DNNs’ non-convexity to obtain over-approximate sound lower and upper bounding functions on $f$, respectively, denoted by $\underline{f}$ and $\overline{f},$ i.e., $\underline{f}({\bm{x}})\leq f({\bm{x}})\leq\overline{f}({\bm{x}})$ for all ${\bm{x}}\in\mathcal{C}.$ Therefore, if $\underline{f}^{*}=\min_{{\bm{x}}\in\mathcal{C}}\underline{f}({\bm{x}})>0$, then also $f^{*}=\min_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})>0$, i.e., the real minimum value of $f$ will be positive, and similarly if $\overline{f}^{*}=\min_{{\bm{x}}\in\mathcal{C}}\overline{f}({\bm{x}})\leq 0$ than also $f^{*}=\min_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})\leq 0$.
In both these situations, we can return a provable result. However, if $\underline{f}^{*}<0<\overline{f}^{*}$, we cannot be sure about the sign of $f^{*}$, and we typically have to proceed with a branch and bound (BaB) (bab) process. More specifically, many FV tools firstly recursively divide the original verification problem into smaller subdomains, either by dividing the perturbation region (reluval) or by splitting ReLU neurons into positive/negative linear domains (babSplitRelu). Secondly, they bound each subdomain with specialized (incomplete) verifiers, typically linear programming (LP) solvers (LP1), which can fully encode neuron split constraints. The verification process ends once we verify all the subdomains of this searching tree, or we find a single counterexample $\bm{x}$} such that $\overline{f}({\bm{x}})\leq 0$. Even though LP-verifiers are mainly used in complete FV tools, recent LiRPA-based approaches (acrown; bcrown) show how to solve an optimization problem that is equivalent to the costly LP-based methods with neuron split constraints while maintaining the efficiency of bound propagation techniques, significantly outperforming LP-verification time thanks to GPU acceleration.
LiRPA approaches (crown; singh2019abstract; autolirpa; acrown; bcrown) propose to cope with the non-linearity of the function computed by a neural network by computing linear approximations of each non-linear unit. The high-level idea is to compute bounds on each neuron’s function in the DNN, for instance, all the ReLU nodes, that can be expressed by linear functions, for which the above robustness verification problem can be efficiently solved. In detail, using interval bound propagation (IBP) (lomuscio2017approach), we first compute a reachable set for each neuron $z^{(i)}_{j}$.
Definition 2.2 (Reachable set).
Given an input perturbation region $\mathcal{C}\subseteq\mathbb{R}^{d}$, the reachable set of a neuron $z^{(i)}_{j}$ is defined as the interval $l^{(i)}_{j},u^{(i)}_{j}$], where $l^{(i)}_{j}\leq z^{(i)}_{j}(\mathbf{x})\leq u^{(i)}_{j}\quad\text{for all }\mathbf{x}\in\mathcal{C}$.
@@ 2.3 Linear Relaxation-based Perturbation Analysis (LiRPA) Approaches
That is, $l^{(i)}_{j}$ and $u^{(i)}_{j}$ represent lower and upper bounds, respectively, on the pre-activation values of neuron $z^{(i)}_{j}$ over the entire perturbation region. The ReLU post-activation value of the node is given by $\hat{z}^{(i)}_{j}=\max(0,z^{(i)}_{j})$. Therefore, the node is considered “unstable" if its pre-activated bounds are such that $u^{(i)}_{j}>0>l^{(i)}_{j}$ and a linear approximate bound can be computed as depicted in Fig. 1. Once linear bounds are established across all neurons, two propagation methods are typically employed: forward and backward. In forward propagation, the linear bounds for each neuron are expressed in terms of the input and propagated layer by layer until the output is reached. In backward propagation, we start from the output and propagate the bounds backward to earlier layers until we can express a linear relation between input and output.
Algorithm 1 LiRPAcrown backward output bounds computation
1:Input: A DNN $f$ with NN layers, and input ${\bm{x}}_{0}$ and an $\varepsilon$ perturbation to compute the perturbation region $\mathcal{C}$, interm_bounds (optional).
2:Output: provable lower bound of $f$ when considering $\mathcal{C}$.
3:if interm_bounds $\emptyset$ then
4:  interm_bounds $\leftarrow\texttt{IBP}(f,\mathcal{C})$
[…]
6:for $i\in\{N-2,\ldots,1\}$ do
[…]
12:return lower_bound
@@
Otherwise, the node is either considered “active" if $l^{(i)}_{j}\geq 0$ or “inactive" if $u^{(i)}_{j}\leq 0$. In detail, to improve the tightness of the bounds, (autolirpa) proposes a refined backward computation strategy that leverages information from a prior forward propagation. We report in Alg. 1 a brief overview of the method. After computing all reachable sets using a method such as interval bound propagation lomuscio2017approach, the backward analysis constructs layer-wise linear relaxations, starting from the output layer and working backward to the input. For clarity purposes, in the following, we will only report the notation for the linear lower bound computation, but similar considerations also apply to the upper bound.